A stack of playing cards contains 4 jacks, 5 queens, 3 kings, and 3 aces. Two cards will be randomly selected from the stack. What is the probability that a queen is chosen and replaced, and then a queen is chosen again?
A.1/9
B.4/9
C.16/225
D.2/21

Answer by rothauserc(4718)

there are 15 cards in the deck
:
Probability of choosing a queen is 5/15 = 1/3
:
Probability of choosing 2 queens with replacement is (1/3) * (1/3) = 1/9
:
answer is A
: